# pnasresearcharticle

### v6

@@ Appendix B Theoretical Analysis
Computation of p-values: We employed the dynamic programming approach described by Noble, et al. (32) to compute the p-values associated with the XCorr scores. This method estimates the distribution of XCorr scores under the null hypothesis and determines the statistical significance of the observed XCorr scores.
@@
Theoretically, the p-value is inversely related to peptide recall. This inverse relationship arises because, when computing the Sidak-corrected p-value (38), we must adjust for a significantly larger number of peptide hypotheses in de novo sequencing than in database searches. In database searches, only the peptides existing in the database are considered, limiting the number of comparisons. In contrast, de novo sequencing must account for all possible peptide sequences, dramatically increasing the number of hypotheses and necessitating a stricter correction for multiple testing. Consequently, the p-value threshold becomes more stringent in de novo sequencing, potentially reducing peptide recall. The relationship between peptide recall and p-value is given by the following equation:
@@ Appendix D Ion Offsets
List of supportingtables 1 Ion offsets for ion types we used. [N] is the molecular mass of the neutral N-terminal group; [C] is the molecular mass of the neutral C-terminal group. C,H,O are the mass of the carbon atom, hydrogen atom and oxygen atom individually. (14)
@@ Appendix F Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA (20). In our experiments, we implemented the detection results from DIA-NN (27) during the main experiment, and DIAUmpire (43) during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) (36), Taynova et al. (2016) (44) or Tsou et al. (2015) (43). The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix H Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 (27) with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 (43).
@@ Appendix I Model Implementation Details
Our model is trained on a Lion optimizer (47) with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
@@
In this ablation study, illustrated in Figure 2, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The removal of coelution-aware pretraining reduces peptide recall by 8.5% compared to DIANovo. The removal of Rotary Positional Encoding (RoPE) results in peptide recall decreasing to 67.2% of the original. Excluding extended peak features or limiting the input to only three neighboring spectra leads to moderate decreases in performance. Finally, binning spectra into 0.5 Th intervals produces a substantial reduction in performance.
@@
We generate a t-SNE (48) plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 3, 4, and 5) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 3, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.

### v7

@@ Appendix B Theoretical Analysis
Computation of p-values: We employed the dynamic programming approach described by Noble, et al. [32] to compute the p-values associated with the XCorr scores. This method estimates the distribution of XCorr scores under the null hypothesis and determines the statistical significance of the observed XCorr scores.
@@
Theoretically, the p-value is inversely related to peptide recall. This inverse relationship arises because, when computing the Sidak-corrected p-value [38], we must adjust for a significantly larger number of peptide hypotheses in de novo sequencing than in database searches. In database searches, only the peptides existing in the database are considered, limiting the number of comparisons. In contrast, de novo sequencing must account for all possible peptide sequences, dramatically increasing the number of hypotheses and necessitating a stricter correction for multiple testing. Consequently, the p-value threshold becomes more stringent in de novo sequencing, potentially reducing peptide recall. The relationship between peptide recall and p-value is given by the following equation:
@@ Appendix D Ion Offsets
List of supportingtables 1 Ion offsets for ion types we used. [N] is the molecular mass of the neutral N-terminal group; [C] is the molecular mass of the neutral C-terminal group. C,H,O are the mass of the carbon atom, hydrogen atom and oxygen atom individually. [14]
@@ Appendix F Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA [20]. In our experiments, we implemented the detection results from DIA-NN [27] during the main experiment, and DIAUmpire [43] during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) [36], Taynova et al. (2016) [44] or Tsou et al. (2015) [43]. The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
@@ Appendix H Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 [27] with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 [43].
@@ Appendix I Model Implementation Details
Our model is trained on a Lion optimizer [47] with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
@@
In this ablation study, illustrated in Supporting Figure 2, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The removal of coelution-aware pretraining reduces peptide recall by 8.5% compared to DIANovo. The removal of Rotary Positional Encoding (RoPE) results in peptide recall decreasing to 67.2% of the original. Excluding extended peak features or limiting the input to only three neighboring spectra leads to moderate decreases in performance. Finally, binning spectra into 0.5 Th intervals produces a substantial reduction in performance.
@@
We generate a t-SNE [48] plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 3, 4, and 5) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 3, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.